100kpimpcess:

ok so we have to use the distributive property so this is the question: 2x-4(x-2)=-7+2x-1

Stuartbent:

I can tell u the answer but ion know how to explain

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Stuartbent I can tell u the answer but ion know how to explain \(\color{#0cbb34}{\text{End of Quote}}\) ok

Stuartbent:

So value of x is 4

100kpimpcess:

ok

umm:

Given\[2x−4(x−2)=−7+2x−1\]Simplify both sides,\[2x−4(x−2)=−7+2x−1\]Distribute\[2x+(−4)(x)+(−4)(−2)=−7+2x+−1\]\[2x+−4x+8=−7+2x+−1\]Combine Like Terms\[(2x+−4x)+(8)=(2x)+(−7+−1)\]\[−2x+8=2x+−8\]\[−2x+8=2x−8\]Subtract 2x from both sides\[−2x+8−2x=2x−8−2x\]\[−4x+8=−8\]Subtract 8 from both sides\[−4x+8−8=−8−8\]\[−4x=−16\]Divide both sides by -4\[\frac{ -4x }{-4}=\frac{ -16 }{ -4 }\]\[x=4\]Therefore, your answer is \[x=4\]

Stuartbent:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @umm Given\[2x−4(x−2)=−7+2x−1\]Simplify both sides,\[2x−4(x−2)=−7+2x−1\]Distribute\[2x+(−4)(x)+(−4)(−2)=−7+2x+−1\]\[2x+−4x+8=−7+2x+−1\]Combine Like Terms\[(2x+−4x)+(8)=(2x)+(−7+−1)\]\[−2x+8=2x+−8\]\[−2x+8=2x−8\]Subtract 2x from both sides\[−2x+8−2x=2x−8−2x\]\[−4x+8=−8\]Subtract 8 from both sides\[−4x+8−8=−8−8\]\[−4x=−16\]Divide both sides by -4\[\frac{ -4x }{-4}=\frac{ -16 }{ -4 }\]\[x=4\]Therefore, your answer is \[x=4\] \(\color{#0cbb34}{\text{End of Quote}}\) This is direct answerr I'm sorry